1.Calculate mass of 2.5 gram atoms of magnesium 2. Calculate the mass of 0.72 gram molecule of CO2 In first question we are multiplying whereas in 2nd we are dividing why?? Why we are not multiplying same in 2nd

Grade:12th pass

1 Answers

shreya
5 years ago
Mass of 1 gram atom of mg=24g
So mass of 2.5 g atom of mg =2.5*24=60g
In this we are multiplying because we have to find of more than1 g so we are multiplying
In second 
Molecular mass of co2 in grams =44g
Mass of 0.72molecular of co2=44*0.72=31.68g
We are not dividing although answer is less than mass of co2 because we are taking out0.72 g from it .........hope it helps 
Thank you....